OpenStudy (anonymous):

i will go answer other qs while u do this. Evaluate the expression 7x2y, when x = 5 and y = 3. 210 315 525 728

OpenStudy (anonymous):

the 2 is a exponent

OpenStudy (anonymous):

not a real 2

OpenStudy (anonymous):

7(5)^2(3)=7*25*3=525

OpenStudy (anonymous):

wow you really are a human calculator.

syed98 (syedmohammed98):

7(5)^2 * 3 175*3 525

OpenStudy (anonymous):

thank you i will fan you.

syed98 (syedmohammed98):

ur welcome :)

OpenStudy (anonymous):

srry this is closed and i frogot to do that